# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

### v1

@@ 7. User Study
  • RQ3 (Performance): What is the perceived satisfaction with PrototypeAgent compared to existing image generation tools?
  • RQ4 (Usage): What is the perceived usefulness of our PrototypeAgent compared to the state-of-the-art industrial tools in terms of the five findings identified in Section 3?
7.0.1. Participants
For both user studies, We recruited the participants through alumni network and partnerships. As a result, we obtained 16 UI/UX practitioners (6 males and 10 females) from various corporate entities to attend our user study. The working experience of these professionals varies, allowing for a broad insight into industry practices and expectations. We had 8 participants with 1-3 years of experience, 5 with 3-5 years, and a further 3 who have been active in the field for more than 5 years. These practitioners come from a spectrum of roles within the UI/UX domain, including UI/UX Designers, UX Engineers, Interaction Designers, and UX Researchers.
@@ 7.1.1. Procedure
Building on established GUI and image evaluation methods (Reinecke et al., 2013; Zhang et al., 2008), participants assessed the quality of mobile GUI designs based on three metrics: functional semantics, design aesthetics, and color harmony. Functional semantics evaluates the relevance and clarity of the generated content, focusing on how well it aligns with the intended functions and the quality of its meaning. Design aesthetics assessed visual appeal (Zhao et al., 2021), and color harmony examined the effectiveness of color combinations. Each design was rated on a 5-point Likert scale. For practical evaluation, 20 design tasks from the Screen2Word dataset (Wang et al., 2021) were selected, and three prototypes were generated per task using both PrototypeAgent and stable diffusion models for comparison. Participants were briefed on these metrics before evaluating the generated GUI designs. They independently scored each design, with the source model of each design concealed to maintain objectivity.
@@ 7.1.2. Results & Discussion
Furthermore, we found that most of our results demonstrated accurate semantic alignment with the design intent, due to our decoupled generation mechanism. This was evident from the high score of 4.32. Comparatively, this score was significantly higher than that of the stable-diffusion-1-5 and stable-diffusion-2-1 generated results 1.58 and 1.7, yet just 0.19 score lower than the real screenshots (4.51). As demonstrated in Figure 7, the stable-diffusion models largely presented images lacking specific semantic content, thus appearing blurry.
Analysis of Design Aesthetics and Color Harmony. Interestingly, a few of the prototypes generated by our model scored higher than real-world prototypes - a notable accomplishment when compared against the benchmark of real screenshots. In Figure 8 (b), our model surpassed the average score for real prototypes in terms of design aesthetics (3.83 compared to 3.71) and in Figure 8 (c) was slightly lower with respect to colour harmony (3.62 compared to 3.74). This is a significant improvement over the stable-diffusion models. Our generated prototypes closely mirrored real-world prototypes’ overall aesthetics and colour harmony and, in some instances, were of superior quality to poorly designed real-world prototypes.
@@
Answer to RQ3: The user study underscored the efficacy of PrototypeAgent in creating GUI designs that excel in design-prototype consistency and aesthetic appeal and colour harmony compared to the state-of-the-art image generation models. Notably, the outputs were found to align closely with real-world screenshots, even outperforming them in terms of design aesthetics. Further analysis underscored that our PrototypeAgent decoupled generation mechanism achieved accurate semantic alignment with the design intent.
Figure 8. Comparative Evaluation of Prototype Generation. “Real” refers to results obtained from corresponding real screenshot images, “s-d-1-5” represents the outcomes of the “stable-diffusion-1-5” model, “s-d-2-1” represents the outcomes of the “stable-diffusion-2-1” model, and “Our” indicates results generated by our PrototypeAgent.
@@ 7.2. User Study 2: Usage Study
In order to obtain insightful feedback on our tool, we compared our PrototypeAgent with emerging design tools; we conducted another user study followed by expert interviews. This study involved a comparative analysis using our tool, Vercel’s V0(Vercel, 2024)—a UI code-based design tool, and Uizard(Uizard, 2024)—a Prompt to UI design tool. After engaging with these tools, participants completed a 5-point Likert Scale questionnaire based on 5 questions. The questionnaire aimed to assess how well our system addressed five design goals identified in Section 3, in terms of all aspects of performance and interactiveness.
@@ 7.2.1. Procedure
    Was the input method intuitive, and did it allow you to express the design intent clearly and precisely? (Related to F2: Interaction Modality)
@@
    Did the tool provide useful suggestions to clarify your design intent? (Related to F3: Implicit Intent Clarification)
@@
    Were the intermediate results useful in aligning the design with your intent, and did they enhance interaction? (Related to F4: Human-Centered Intent-Design Alignment)
@@
    Did the tool maintain consistent semantics and style across all design elements, ensuring a cohesive output? (Related to F5: Ensuring Design Coherence)
Figure 9. Results of the usage study. Blue represents Vercel’s V0, green denotes Uizard, and red indicates our system PrototypeAgent.
@@ 7.2.2. Results & Discussion
This feedback highlights the potential of our tool in integrating design libraries—a feature lacking in Vercel’s V0 and partially implemented in Uizard. Our PrototypeAgent offers a more customizable and functional design experience.
Interaction Modality (F2). Our system led in reflecting design intent with a score of 4.1, compared to Uizard’s 3.6 and Vercel’s 2.5. Participants valued our tool’s use of both wireframes and prompts, which Uizard lacks, and found Vercel’s code generation to not adhere closely to layout sizes. A participant remarked, “PrototypeAgent’s method of rendering components from wireframes and prompts is impressive. The design of layout parts encompasses functional design and is labor-intensive. AI assistance in this area is much needed.”
Implicit Intent Clarification (F3). Our PrototypeAgent scored 4.0 for Implicit Intent Clarification, outperforming Uizard (3.2) and Vercel’s V0 (1.6). 14/16 participants found our clarification ability particularly helpful, especially for themes and app categories. One participant noted, ”The automatic clarification is more professional than my own input, saving me from guessing what the AI needs.” This feature allowed designers to focus more on achieving their desired results, reducing guesswork and effort. On average, designers using our tool made 2.6 revisions to reach their desired design. In comparison, Uizard required 4.3 revisions, and Vercel’s V0 required 6.1 revisions, with many users still unsatisfied, citing lack of control over the generated code format.
@@
Human-Centered Intent-Design Alignment(F4). Our tool was highly regarded for its intermediate generation process in helping with human-in-the-loop intent-design alignment, achieving a score of 3.8. In comparison, Uizard scored 3.5, while Vercel’s V0 lagged behind with a 2.5. Notably, 14 out of 16 participants found the intermediate steps in our tool exceedingly beneficial for interactive engagement with the AI. A participant shared a compelling example: “While working on a job search page, I was impressed by the UI designs produced by the tool. They were professional and included all necessary details, such as job categories, locations, and salaries, making it feel like a real, complete design.” This feedback underscores our tool’s proficiency in providing clear, informative intermediate steps that facilitate effective user interaction and ensure the generation of high-quality prototypes.
@@
Ensuring Design Coherence (F5). In the critical aspect of maintaining design coherence, our tool and Uizard both achieved a high score of 4.2, clearly outperforming Vercel’s V0’s 3.5. A majority of the participants, 10 out of 16, remarked on the effective maintenance of colour and image style consistency by both our tool and Uizard. They noted that these tools produced designs with cohesive colour schemes and consistent visual elements.
@@
Weaknesses and improvements. Participants offered valuable feedback on potential improvements and identified exciting directions for future enhancements. Several participants saw great potential in augmenting the user experience by introducing automated wireframe generation from high-level descriptions. This feature could significantly streamline the design process by reducing manual steps, thereby increasing efficiency. A participant highlighted this by saying, “It would be really helpful if the tool could auto-generate wireframes based on high-level descriptions. It would save me more time.”
@@ 8. Discussion and Future Work
Our current work, PrototypeAgent, focuses on improving the design generation process and enhancing the quality of prototypes. It draws inspiration from a rich knowledge base and primarily relies on high-level design descriptions and wireframe layouts to translate design intent into tangible design prototypes. Looking ahead, there are several future directions for improvements and extensions:
1) Exploring Enhanced AI-Human Collaboration in UI Design A key future direction involves refining input modalities to support a broader range of intent-guided inputs, such as sketching wireframes or referencing design guidelines. The aim is not merely to expand the input options but to minimize user effort while helping designers express their intent more precisely with less input. For instance, our current system’s Implicit Intent Clarification feature already reduces user workload by automatically clarifying and refining design intent.
Additionally, we can explore how AI can further assist designers by automating more aspects of the process, such as generating wireframes automatically, thereby reducing manual effort. Although it would be relatively straightforward to remove the wireframe creation input and allow our system to automatically retrieve and generate wireframes, our current focus is on achieving accurate input expression. For this reason, we chose not to include this feature at this stage. However, we recognize the potential benefits of automated wireframe generation and see it as a valuable direction for future enhancements.
2) Usability and Accessibility Enhancements. As we continue to improve PrototypeAgent, we are committed to augmenting our system with standard mobile device user interface guidelines like material design (Google, 2023). These improvements will further enhance our system’s usability and accessibility, ensuring our prototypes not only look appealing but also offer intuitive and user-friendly interactions. To further improve the design process, we could incorporate usability testing and visual reports that highlight potential issues early in the design stage, even during the generation process. Early detection and correction of design flaws will help reduce the cost of revisions, making the system more efficient and accessible.
3) Contextual Relationships and Multi-page Designs. Recognizing that UI designs often span multiple pages, in the future, we could discuss and expand on contextual relationships in our design generation. This will allow us to produce designs for multiple pages simultaneously in future iterations, ensuring a cohesive style across these pages and thereby enhancing efficiency.
4) Tool Integration and Design-to-Code. With an aim to seamlessly incorporate our system into designers’ workflow, we imagine PrototypeAgent could be developed as a plugin that integrates smoothly with popular wireframe sketching software, such as Figma (Figma, 2024) and Sketch (Sketch, 2024). This would widen our system’s functionality by providing designers with robust editing capabilities directly within these platforms. Additionally, future work could also explore ways to automate design-to-code conversion to provide designers and developers with an end-to-end solution.

### v3

@@ 7. User Study
  • RQ3 (Performance): What is the perceived satisfaction with PrototypeFlow compared to existing image generation tools?
  • RQ4 (Usage): What is the perceived usefulness of our PrototypeFlow compared to the state-of-the-art industrial tools in terms of the five findings identified in Section 3?
  • RQ5: (Case): What are the strengths and weaknesses of different system variants, and how can they inform future design tools?
Participants: For both user studies, We recruited the participants through alumni network and partnerships. As a result, we obtained 16 UI/UX practitioners (6 males and 10 females) from various corporate entities to attend our user study. The working experience of these professionals varies, allowing for a broad insight into industry practices and expectations. We had 8 participants with 1-3 years of experience, 5 with 3-5 years, and a further 3 who have been active in the field for more than 5 years. These practitioners come from a spectrum of roles within the UI/UX domain, including UI/UX Designers, UX Engineers, Interaction Designers, and UX Researchers.
@@ 7.1.1. Procedure
Building on established GUI and image evaluation methods (Reinecke et al., 2013; Zhang et al., 2008), participants assessed the quality of mobile GUI designs based on three metrics: functional semantics, design aesthetics, and color harmony. Functional semantics evaluates the relevance and clarity of the generated content, focusing on how well it aligns with the intended functions and the quality of its meaning. Design aesthetics assessed visual appeal (Zhao et al., 2021), and color harmony examined the effectiveness of color combinations. Each design was rated on a 5-point Likert scale. For practical evaluation, 20 design tasks from the Screen2Word dataset (Wang et al., 2021) were selected, and three prototypes were generated per task using both PrototypeFlow and stable diffusion models for comparison. Participants were briefed on these metrics before evaluating the generated GUI designs. They independently scored each design, with the source model of each design concealed to maintain objectivity.
Figure 9. Generated GUI Examples from Input Layout and Prompt: Corresponding Rico Screenshots, stable-diffusion-1-5, stable-diffusion-2-1, and Our PrototypeFlow.
@@ 7.1.2. Results & Discussion
Furthermore, we found that most of our results demonstrated accurate semantic alignment with the design intent, due to our decoupled generation mechanism. This was evident from the high score of 4.32. Comparatively, this score was significantly higher than that of the stable-diffusion-1-5 and stable-diffusion-2-1 generated results 1.58 and 1.7, yet just 0.19 score lower than the real screenshots (4.51). As demonstrated in Fig. 9, the stable-diffusion models largely presented images lacking specific semantic content, thus appearing blurry.
Analysis of Design Aesthetics and Color Harmony. Interestingly, a few of the prototypes generated by our model scored higher than real-world prototypes - a notable accomplishment when compared against the benchmark of real screenshots. In Fig. 10 (b), our model surpassed the average score for real prototypes in terms of design aesthetics (3.83 compared to 3.71) and in Fig. 10 (c) was slightly lower with respect to colour harmony (3.62 compared to 3.74). This is a significant improvement over the stable-diffusion models. Our generated prototypes closely mirrored real-world prototypes’ overall aesthetics and colour harmony and, in some instances, were of superior quality to poorly designed real-world prototypes.
@@
Answer to RQ3: The user study underscored the efficacy of PrototypeFlow in creating GUI designs that excel in design-prototype consistency and aesthetic appeal and colour harmony compared to the state-of-the-art image generation models. Notably, the outputs were found to align closely with real-world screenshots, even outperforming them in terms of design aesthetics. Further analysis underscored that our PrototypeFlow decoupled generation mechanism achieved accurate semantic alignment with the design intent.
Figure 10. Comparative Evaluation of Prototype Generation. “Real” refers to results obtained from corresponding real screenshot images, “s-d-1-5” represents the outcomes of the “stable-diffusion-1-5” model, “s-d-2-1” represents the outcomes of the “stable-diffusion-2-1” model, and “Our” indicates results generated by our PrototypeFlow.
@@ 7.2. User Study 2: Usage Study
In order to obtain feedback and facilitate further discussion on our findings and tool usage, we compared our PrototypeFlow with emerging design tools; we conducted another user study followed by expert interviews. This study involved a comparative analysis using our tool, Vercel’s V0(Vercel, 2024)—a UI code-based design tool, and Uizard(Uizard, 2024)—a Prompt to UI design tool. After engaging with these tools, participants completed a 5-point Likert Scale questionnaire based on 5 questions. The questionnaire aimed to assess how well our system addressed five design goals identified in Section 3, in terms of all aspects of performance and interactiveness.
@@ 7.2.1. Procedure
    Was the input method intuitive, and did it allow you to express the design intent clearly and precisely? (Related to F2: More Input Control and Flexible Output Editability in Design Generation Process)
@@
    Did the tool provide useful suggestions to clarify your design intent? (Related to F3: Supporting Designers in Expressing Intent Through Prompts)
@@
    Were the intermediate results useful in aligning the design with your intent, and did they enhance interaction? (Related to F4: Precise Control in Generation Processes)
@@
    Did the tool maintain consistent semantics and style across all design elements, ensuring a cohesive output? (Related to F5: Maintaining Thematic Consistency and Coherence Across Generated Components)
Figure 11. Results of the usage study. Blue represents Vercel’s V0, green denotes Uizard, and red indicates our system PrototypeFlow.
@@ 7.2.2. Results & Discussion
For statistical analysis, we conducted independent-samples t-tests comparing PrototypeFlow with Vercel’s V0 and Uizard across five user survey question criteria (F1–F5). Results show that PrototypeFlow significantly outperformed Vercel on all dimensions (p ¡ 0.001). Compared to Uizard, PrototypeFlow showed significant advantages in intermediate transparent and editable generation (F4), suggestion usefulness in clarifying and enhancing designer prompts (F3), and input expressiveness (F2), all with p ¡ 0.001. For design theme consistency (F5), PrototypeFlow and Uizard performed similarly (p = 0.071), and for relevance of retrieved suggestions (F1), PrototypeFlow scored slightly lower (3.5 vs. 3.6, p = 0.4). These findings indicate that PrototypeFlow excels at enhancing clarity, interaction, and refinement in the design process, while maintaining comparable output consistency.
@@
This feedback highlights the potential of our tool in integrating design libraries—a feature lacking in Vercel’s V0 and partially implemented in Uizard. Our PrototypeFlow offers a more customizable and functional design experience.
More Input Control and Flexible Output Editability in Design Generation Process (F2). Our system led in reflecting design intent with a score of 4.1, compared to Uizard’s 3.6 and Vercel’s 2.5. Participants valued our tool’s use of both wireframes and prompts, which Uizard lacks, and found Vercel’s code generation to not adhere closely to layout sizes. A participant remarked, “PrototypeFlow’s method of rendering components from wireframes and prompts is impressive. The design of layout parts encompasses functional design and is labor-intensive. AI assistance in this area is much needed.”
Supporting Designers in Expressing Intent Through Prompts (F3). Our PrototypeFlow scored 4.0 for Prompt Enhancement, outperforming Uizard (3.2) and Vercel’s V0 (1.6). 14/16 participants found our clarification ability particularly helpful, especially for themes and app categories. One participant noted, ”The automatic clarification is more professional than my own input, saving me from guessing what the AI needs.” This feature allowed designers to focus more on achieving their desired results, reducing guesswork and effort. On average, designers using our tool made 2.6 revisions to reach their desired design. In comparison, Uizard required 4.3 revisions, and Vercel’s V0 required 6.1 revisions, with many users still unsatisfied, citing lack of control over the generated code format.
@@
Precise Control in Generation Processes (F4). Our tool was highly regarded for its intermediate generation process in helping with human-in-the-loop interaction, achieving a score of 3.8. In comparison, Uizard scored 3.5, while Vercel’s V0 lagged behind with a 2.5. Notably, 14 out of 16 participants found the intermediate steps in our tool exceedingly beneficial for interactive engagement with the AI. A participant shared a compelling example: “While working on a job search page, I was impressed by the UI designs produced by the tool. They were professional and included all necessary details, such as job categories, locations, and salaries, making it feel like a real, complete design.” This feedback underscores our tool’s proficiency in providing clear, informative intermediate steps that facilitate effective user interaction and ensure the generation of high-quality prototypes.
@@
Maintaining Thematic Consistency and Coherence Across Generated Components (F5). In the critical aspect of maintaining design coherence, our tool and Uizard both achieved a high score of 4.2, clearly outperforming Vercel’s V0’s 3.5. A majority of the participants, 10 out of 16, remarked on the effective maintenance of colour and image style consistency by both our tool and Uizard. They noted that these tools produced designs with cohesive colour schemes and consistent visual elements.
@@
7.2.3. Weaknesses and improvements.
Participants offered valuable feedback on potential improvements and identified exciting directions for future enhancements. Several participants saw great potential in augmenting the user experience by introducing automated wireframe generation from high-level descriptions. This feature could significantly streamline the design process by reducing manual steps, thereby increasing efficiency. A participant highlighted this by saying, “It would be really helpful if the tool could auto-generate wireframes based on high-level descriptions. It would save me more time.”
@@
7.3. User Study 3: Case Study
To explore greater flexibility and understand the trade-off between precise control and design flexibility, we introduce a case study that envisions idealized variations of our system and examines their strengths, limitations, and potential directions for future development.
7.3.1. Experimental Setup.
We explore four types of variations across both input and output dimensions. For input variants, we consider (1) Input-NoLayout – removing the layout constraint to allow generation of UIs that preserve semantic structure without adhering to a fixed layout; (2) Input-NoKnowledge – removing the knowledge reference to enable the generation of more diverse and unconstrained applications. For output variants, we consider (3) Output-MultiLayout – generating multiple UI prototypes with varying layouts for the same semantic and thematic specification; (4) Output-MultiTheme – generating UI prototypes that maintain consistent layout and semantics but vary in thematic styles.
To ground these variations, we apply our system to a scenario of designing a self-help search page for a Google product. For each variant, we construct a representative and idealized output that reflects the relaxed constraint as shown in Fig. 12. These outputs serve as design probes to support structured reflection. We then re-engaged the same 16 designers from our previous study and invited them to evaluate each variation and compare. They were asked to reflect on the perceived value, applicable design stage, and trade-offs of each alternative. After obtaining feedback from the designers, we conducted a thematic analysis of the interview data. All responses were transcribed and open-coded to identify patterns in designers’ perceptions of each variant’s usefulness, strengths, and limitations. Two researchers independently coded a subset of the data and resolved differences through discussion to ensure consistency.
7.3.2. Findings.
After analyzing the feedback and the open-coding methods, we derived 3 key findings:
Input Preferences During Ideation: 13 out of 15 designers—especially those with over three years of experience—expressed a preference for the Input–NoKnowledge variant during the early ideation stage. They felt that removing the knowledge constraint allowed for broader and more diverse design suggestions, which helped stimulate creative exploration. These designers noted that relying solely on the internal knowledge base at this stage tended to limit novelty and inspiration. By contrast, the Input–NoLayout variant was perceived as less beneficial during ideation, as removing layout constraints often resulted in outputs that were too detached from practical or implementable UI structures.
Wireframes as a Preferred Input Modality: Designers consistently favored wireframes as their primary method of interaction, emphasizing that this input format aligns with their established workflows and offers precise control over layout and structure. Many described wireframes as a “native language” for communicating design intent—more intuitive and efficient than crafting detailed textual prompts. Several participants noted they maintained personal libraries of reusable wireframes, enabling rapid adaptation to new design tasks. This strong preference reflects a desire to retain detailed layout control by default, and only strategically apply variants like Input–NoLayout when seeking greater output diversity.
Output Diversity Preferences: Helpful for Learning, Less Aligned with Team Workflows. The Output–MultiLayout and Output–MultiTheme variants—representing post-generation diversity—were appreciated primarily by less-experienced designers (4 out of 15, with fewer than two years of experience). For these participants, generating multiple design alternatives after the initial prototype served as a valuable learning tool, enabling comparison, self-validation, and a deeper understanding of design trade-offs. However, even these designers noted that such diversity is less applicable in real-world team settings, where key layout and functionality decisions are typically made collaboratively before generation begins. Most experienced designers (11 out of 15) shared this view, emphasizing that once a team reaches alignment, generating additional alternatives can distract from execution and reduce efficiency. They preferred to front-load diversity during early ideation (e.g., through Input–NoKnowledge), and viewed post-generation variation as potentially disruptive to streamlined workflows and delivery timelines.
Figure 12. Illustration of ideal outputs from different variants of our PrototypeFlow.
Answer to RQ5: Our findings indicate that the timing and form of diversity introduction should align with the designer’s experience and workflow context. Input-level variants like Input–NoKnowledge were preferred by experienced designers during early ideation, enabling broader exploration. In contrast, output-level variants such as Output–MultiLayout and Output–MultiTheme were more helpful for novice designers to compare alternatives and learn from variation. However, these output variants were seen as less useful in team-based settings, where design decisions are often finalized early. These results suggest that future systems should allow dynamic adjustment between control and flexibility, tailored to users’ experience levels and design stages.
@@ 8. Discussion and Future work
The rapid rise of generative models and prompt engineering is transforming many landscapes, including design tools. Our findings highlight the need for GenAI design systems that balance fine-grained designer control with diverse output and adapt to a broad spectrum of users and design goals. Managing and updating design knowledge bases remains a challenge, as static resources can quickly become outdated in fast-evolving contexts. This section discusses key implications from our system design and user studies and identifies future directions, including broader user and platform support, dynamic knowledge management, and enhanced adaptability to emerging design trends.
8.1. Tailoring Interaction Modalities to User Expertise
Our research primarily targeted professional UI/UX designers and found that these experts prefer to interact with AI design tools in their own “native” design languages, primarily wireframes and visual layouts. They noted that this approach gives them precise control over structure and aligns with their established workflows. One interviewee described wireframes as the “native language” for communicating design intent, more intuitive than lengthy text prompts. This insight underscores a key design principle: AI tools should “align with designers’ visual habits” and not force them into unfamiliar interaction styles(Khan et al., 2025). In practice, this means integrating GenAI assistance into existing GUI design environments (e.g. Figma, Adobe XD) where designers naturally work with visual components, rather than requiring designers to write long textual descriptions of UIs. Indeed, recent studies have found that designers want AI helpers to “go beyond textual prompting” and seamlessly fit into visual ideation processes(Khan et al., 2025). Our PrototypeFlow reflects this by letting designers sketch a layout or provide a wireframe which the system then fleshes out, rather than relying on abstract prompts alone.
However, not all users of AI-assisted design tools will be professionals. Novice designers or non-designers, such as developers or small entrepreneurs, may lack the training to sketch wireframes or use design terminology. For these users, the system must “speak” a different language – one that matches their mental model and skills. Prior work has identified this gap: only focusing on experienced designers leaves out the perspectives of novice designers and students, whose needs and ways of communicating with AI might differ significantly (Khan et al., 2025). For example, a beginner might prefer describing the intended interface in plain language (e.g. “I need an e-commerce home page with a search bar at the top and product cards”) or selecting from example images or templates, rather than drawing a layout from scratch, as they don’t have their own wireframe design library. There is potential for GenAI design tools to incorporate multiple input modalities to accommodate this range of users. A system could allow a user to start from a high-level natural-language description or a storyboard of screenshots, and then iteratively refine the design with more visual inputs as their confidence grows. In essence, the AI should serve as a fluent translator between different “UI languages” – whether that’s the precise vocabulary of a UX professional or the rough descriptions a novice might give. By tailoring the interaction style (textual, visual, conversational, or guidance-based) to the user’s expertise, we can lower the barrier to entry and democratize the design process. Our findings that experts prefer wireframe-driven interaction suggests one size won’t fit all; future systems might include an adaptive interface that, for instance, starts with a friendly Q&A or template gallery for novices and gradually introduces more free-form wireframing as the user gains proficiency. Such adaptability would help non-professionals work in whatever way feels most natural to them, turning AI into a true creative partner rather than a rigid tool.
8.2. Balancing Fine-Grained Control and Output Diversity in Generative Design
A recurring topic is the tension between giving designers fine-grained control over outputs, while providing high output diversity for inspiration. Our findings reflect this balance: senior designers tend to prioritize control and precision, especially in later stages of design (such as creation and iteration), whereas less-experienced designers or early ideation phases benefit from a breadth of diverse suggestions. In our case study, 13 of 15 designers preferred turning off the internal knowledge base during early ideation to avoid constraining the AI to known patterns. This allowed more divergent and novel UI suggestions to emerge. By contrast, turning off layout constraints was seen as less useful at creation – outputs became too unimplementable. This highlights that designers desire structured diversity: room for exploration, but within workable bounds.
Prior work echoes this need for multiple options and iterative exploration; for example, novice spatial designers wished for AI systems to generate multiple design options, so they can choose and refine among them(Wan et al., 2024). At the same time, users criticized fully automated, “one-shot” generation as a black box process that they couldn’t intervene in, preferring approaches (like retrieval-based methods) that allowed more control over intermediate steps(Wan et al., 2024). In the context of UI design, this means an effective GenAI tool should support both divergent exploration and convergent refinement. This aligns with the classic divergent–convergent model of design thinking, the system should encourage divergence (high variety, novelty), and later enable convergence (fine-tuning a chosen direction). Recent HCI research emphasizes balancing these dual needs: designers call for AI tools that help “balance efficiency and exploration”, providing varied design alternatives while still letting them control the process.
Importantly, the value of output diversity appears context-dependent. Our participants noted that generating many alternative layouts or themes after an initial prototype was mainly useful as a learning aid for junior designers (e.g. to compare options and understand trade-offs), but was less aligned with the realities of professional team workflows. Once a design direction is decided collaboratively, excessive variations can become a distraction and slow down execution. Experienced designers in our study preferred to front-load diversity in the early ideation stage to broaden horizons, then commit to a direction and focus on polishing it. This sentiment resonates with findings by Khan et al.(Khan et al., 2025): professional UI/UX designers “valued AI tools that offer greater control over ideation” while still generating design alternatives for inspiration. In sum, an AI design system must intelligently support both modes – offering divergent idea generation when appropriate and convergent fine-tuning when the designer needs precision. Designing interfaces that let users fluidly toggle or transition between these modes (for example, a “surprise me with alternatives” button versus a “lock this layout” mode) could be a promising direction for future tools.
8.3. The Limits of Static Knowledge Bases in Rapidly Evolving Design Contexts
Our PrototypeFlow system features a knowledge-based retrieval module that leverages a local repository of established design patterns and examples (such as company style guidelines) to ground generative outputs in familiar, brand-consistent solutions. This foundation helps ensure results are realistic and on-brand, accelerating routine design tasks with proven patterns. However, our interviews highlight several limitations to this approach, particularly as design contexts and trends rapidly evolve. First, reliance on a static or narrowly scoped knowledge base can quickly lead to outdated or repetitive suggestions, constraining creativity during early ideation. Designers noted in Section 7.3 that while knowledge-based retrieval is helpful for efficiency, over-dependence early in the process risks limiting exploration and novelty. This points to an inherent trade-off between exploiting known best practices and enabling creative divergence. A local or company-specific knowledge base that is too limited or poorly maintained may cause the system to repeatedly generate similar styles, ultimately missing out on new design paradigms and innovative alternatives.
Second, maintaining the relevance and quality of the knowledge base is an ongoing challenge. In real-world team knowledge base deployments, questions arise: Who is responsible for updating the design repository? How are new styles and patterns integrated? If left unattended, the system may propose outdated trends (such as the overuse of gradients or drop shadows), reducing the value of its recommendations. Furthermore, adaptability to diverse design contexts remains an open issue. A knowledge base primarily built from e-commerce app screens may not generalize well to other domains, such as data visualization dashboards or virtual reality interfaces. Flexible retrieval that allows users to constrain or expand the knowledge base according to project needs may help address this, and our system takes a first step by enabling users to toggle knowledge use depending on the task. However, more automation and standardization mechanisms for dynamic updating and domain adaptation are needed.
To address these limitations, future work should explore more dynamic and scalable approaches to knowledge management. For example, integrating external sources such as public design repositories, design galleries, or open-source community platforms (similar to Hugging Face in deep learning) could help ensure broader coverage and more timely updates. Establishing standardized formats for online management and community contributions would further support the evolution and sharing of up-to-date design knowledge.
In summary, while knowledge-based UI generation provides a strong foundation and alignment with established design standards, its long-term effectiveness depends on broad coverage, timely updates, adaptability to new domains, and the ability to learn from ongoing user feedback. Ensuring that GenAI design assistants stay in sync with current trends and remain responsive to diverse project contexts will be key to their sustained usefulness.
8.4. Integrating GUI and Web Design Practices for Broader Design Knowledge
While PrototypeFlow and similar generative AI systems have advanced graphical user interface (GUI) design, particularly for mobile and desktop app screens, it is important to consider how these approaches might extend to web design and other platforms. Designing for the web introduces unique challenges and requires a different mindset from traditional software UI design. As Jakob Nielsen observed(Nielsen, 1997), web designers cannot exercise full control over the user interface, since the end-user’s device, browser, and personal settings all influence the final appearance and behavior. In practice, this means web layouts must be fluid and adaptable to various screen sizes, network conditions, and accessibility requirements, whereas GUI designers for native applications typically operate with more predictable parameters and can specify layouts with greater precision.This fundamental difference highlights a key limitation of current GenAI design tools trained predominantly on mobile app screens or static interfaces. Such systems may not generalize well to web design scenarios, which demand features like responsive layout reflow, hyperlink navigation, and compliance with web standards such as HTML, CSS, and accessibility guidelines. By contrast, GUI design tools often rely on platform-specific conventions, like those prescribed in the iOS Human Interface Guidelines or Android Material Design.
To bridge these domains, a GenAI design system should incorporate context awareness, enabling it to recognize the intended medium and apply design patterns suited to that context. For example, an AI tool generating a website should suggest responsive grid systems and ensure that layouts adapt seamlessly from mobile to desktop. In contrast, for mobile app design, the system might prioritize native components and platform-specific navigation. Prior research in HCI has shown the value of such context sensitivity. For instance, Landay’s DENIM tool (Lin et al., 2001) was developed specifically for early-stage web design, integrating site maps and page flows to address web designers’ needs—needs that were not met by conventional GUI prototyping tools. This example underscores that certain aspects of design knowledge are inherently tied to the target domain.
Despite these differences, some principles—such as visual hierarchy, consistency, and affordances—are universal across both GUI and web design. The next generation of AI design assistants could leverage a broad foundation of design knowledge while dynamically adapting recommendations and outputs to suit the relevant context. Currently, our system focuses on GUI prototyping. To support web design, it would need to accommodate multi-page navigation, fluid layouts, and a wider set of building blocks such as form elements and navigation bars. Achieving true cross-platform capability will require training on diverse datasets, including both mobile app UIs and responsive web designs, as well as offering user options to tailor the generation process to specific platform requirements. By acknowledging and embracing these domain differences, and drawing from expert guidance on design adaptability, future GenAI tools can become more general partners for designers, supporting creative work across an expanding range of digital interfaces.